# Neuro-Symbolic Entity Alignment via Variational Inference

### v1

@@ Abstract
Entity alignment (EA) aims to merge two knowledge graphs (KGs) by identifying equivalent entity pairs. Existing methods can be categorized into symbolic and neural models. Symbolic models, while precise, struggle with substructure heterogeneity and sparsity, whereas neural models, although effective, generally lack interpretability and cannot handle uncertainty. We propose NeuSymEA, a probabilistic neuro-symbolic framework that combines the strengths of both methods. NeuSymEA models the joint probability of all possible pairs’ truth scores in a Markov random field, regulated by a set of rules, and optimizes it with the variational EM algorithm. In the E-step, a neural model parameterizes the truth score distributions and infers missing alignments. In the M-step, the rule weights are updated based on the observed and inferred alignments. To facilitate interpretability, we further design a path-ranking-based explainer upon this framework that generates supporting rules for the inferred alignments. Experiments on benchmarks demonstrate that NeuSymEA not only significantly outperforms baselines in terms of effectiveness and robustness, but also provides interpretable results.
@@ 1 Introduction
Knowledge graphs (KGs) are crucial for organizing structured knowledge about entities and their relationships, enhancing search capabilities across various applications. They are widely used in question-answering systems (Zhang et al., 2023a; Dong et al., 2023), recommendation systems (Catherine & Cohen, 2016), social network analysis (Tang et al., 2008), Natural Language Processing (Weikum & Theobald, 2010), etc.. Despite their utility, real-world KGs often face issues like incompleteness, domain specificity, and language constraints, which hinder their effectiveness in cross-disciplinary or multilingual contexts. To address these issues, entity alignment (EA) aims to merge disparate KGs into a unified, comprehensive knowledge base by identifying and linking equivalent entities across different KGs. For example, aligning entities between a biomedical KG and a pharmaceutical KG allows for mining cross-discipline relationships through the aligned entities, such as identifying the same drugs and their effects on different diseases to enhance drug repurposing efforts. This alignment enables more nuanced exploration and interrogation of interconnected data, providing richer insights into how entities function across multiple domains.
@@
To perform entity alignment, existing methods can be broadly categorized into symbolic models and neural models. Symbolic models (Suchanek et al., 2012; Jiménez-Ruiz & Cuenca Grau, 2011; Qi et al., 2021) provide interpretable and precise inference by mining ground rules, but they struggle with aligning long-tail entities, especially those without aligned neighbors. In such cases, the lack of supporting rules leads to low recall. Conversely, neural models, such as translation models (Chen et al., 2017; Sun et al., 2018) and Graph Convolutional Networks (GCNs) (Mao et al., 2020; 2021; Wang et al., 2018; Mao et al., 2022; Li et al., 2024b; a), excel in recalling similar entities by embedding them in a continuous space, yet they often fail to distinguish entities with similar representations, causing a drop in precision as the entity pool grows. Neuro-symbolic models aim to combine the strengths of both approaches, offering the interpretability and precision of symbolic models alongside the high recall capabilities of neural models.
@@
However, neuro-symbolic reasoning in entity alignment (EA) faces several challenges. First, combining symbolic and neural models into a unified framework is suboptimal due to the difficulty in aligning their objectives. Current approaches either use neural models as auxiliary modules for symbolic models to measure entity similarity (Qi et al., 2021) or employ symbolic models to refine pseudo-labels (Liu et al., 2022; Chen et al., 2024b). Second, in EA task, the search space for rules is large, as the EA task requires deriving ground rules from both intra-KG and inter-KG structural patterns, leading to an exponentially large search space with increasing rule length. Finally, generating interpretations for EA remains underexplored. Effective interpretations should not only generate supporting rules but also quantify their confidence through rule weights.
@@
To overcome these challenges, we propose NeuSymEA, a neuro-symbolic framework that combines the strengths of both symbolic and neural models. NeuSymEA models the joint probability of truth score assignment for all possible entity pairs using a Markov random field, regulated by a set of weighted rules. This joint probability is optimized via a variational EM algorithm. During the E-step, a neural model parameterizes the truth scores and infers the missing alignments. In the M-step, the rule weights are updated based on both observed and inferred alignments. To leverage long rules without suffering from the exponential search space, we employ logic deduction to decompose rules of any length into shorter, unit-length rules. This allows for efficient inference and weight updates for long rules. After training, the learned rules are adapted into an explainer, enhancing interpretability. Specifically, we reverse the logic deduction process to calculate the weights of long rules based on the learned weights of their shorter counterparts. Additionally, an explainer is introduced to efficiently generate supporting rules with quantified confidence through breadth-first search, as explanations for inferred alignments. Our contributions are summarized as below:
  • A principled neuro-symbolic reasoning framework: NeuSymEA seamlessly integrates neural and symbolic models through a principled variational EM framework, combining their strengths for effective entity alignment.
@@
  • Empirical validation and superior results: NeuSymEA demonstrates state-of-the-art performance on benchmark datasets, delivering both robust alignment accuracy and insightful rule-based interpretations.
@@ 2.1 Problem statement
A knowledge graph ${\mathcal{G}}$ comprises a set of entities ${\mathcal{E}}$, a set of relations ${\mathcal{R}}$, and a set of relation triples ${\mathcal{T}}$ where each triple $(e_{i},r_{k},e_{j})\in{\mathcal{T}}$ represents a directional relationship between its head entity and tail entity. Given two KGs ${\mathcal{G}}=\{{\mathcal{E}},{\mathcal{R}},{\mathcal{T}}\}$, ${\mathcal{G}}^{\prime}=\{{\mathcal{E}}^{\prime},{\mathcal{R}}^{\prime},{\mathcal{T}}^{\prime}\}$, and a set of observed aligned entity pairs ${\mathcal{O}}=\{(e_{i},e_{i}^{\prime})|e_{i}\in{\mathcal{E}},e^{\prime}_{i}\in{\mathcal{E}}^{\prime}\}_{i=1}^{n}$, the goal of entity alignment is to infer the missing alignments by reasoning with the existing alignments. This problem can be formulated in a probabilistic way: each pair $(e,e^{\prime}),e\in{\mathcal{E}},e^{\prime}\in{\mathcal{E}}^{\prime}$ is associated with a binary indicator variable ${\bm{v}}_{(e,e^{\prime})}$. ${\bm{v}}_{(e,e^{\prime})}=1$ means $(e,e^{\prime})$ is an aligned pair, and ${\bm{v}}_{(e,e^{\prime})}=0$ otherwise. Given some observed alignments ${\bm{v}}_{O}=\{{\bm{v}}_{(e,e^{\prime})}=1\}_{(e,e^{\prime})\in{\mathcal{O}}}$, we aim to predict the labels of the remaining hidden entity pairs =×\${\mathcal{H}}={\mathcal{E}}\times{\mathcal{E}}^{\prime}\backslash{\mathcal{O}}$, i.e., ${\bm{v}}_{H}=\{{\bm{v}}_{(e,e^{\prime})}\}_{(e,e^{\prime})\in{\mathcal{H}}}$.
@@ 2.2 Symbolic reasoning for entity alignment
$w_{p,p^{\prime}}:\quad(e_{j}\equiv e_{j}^{\prime})\wedge p(e_{i},e_{j})\wedge p^{\prime}(e_{i}^{\prime},e_{j}^{\prime})\Longrightarrow(e_{i}\equiv e_{i}^{\prime}),$ (1)
@@ 3 Neuro-symbolic reasoning framework for entity alignment
Figure 1: Framework illustration of NeuSymEA. The yellow solid line represents the alignment of an anchor pair. The symbolic model computes the matching probability of entity pairs by mining supporting rules (path pairs from anchor pairs) and evaluating their corresponding weights. The neural model learns embeddings and calculates entity-level matching scores based on embedding similarity. NeuSymEA models the agreement between the symbolic reasoning and neural representations using a joint probability distribution over observed pairs and parameterized truth scores for hidden pairs, optimized through a variational EM algorithm.
@@ 3.1 Variational EM
here, $q({\bm{v}}_{H})$ is a variational distribution of the hidden variables ${\bm{v}}_{H}$. This inequality holds for all $q$ because $p_{w}({\bm{v}}_{O})=\text{ELBO}(q,{\bm{v}}_{O};w)+D_{KL}(q({\bm{v}}_{H})\|p_{w}({\bm{v}}_{H}\mid{\bm{v}}_{O}))$, where $D_{KL}(q({\bm{v}}_{H})\|p_{w}({\bm{v}}_{H}\mid{\bm{v}}_{O}))\geq 0$ is the KL-divergence between $q({\bm{v}}_{H})$ and $p_{w}({\bm{v}}_{H}\mid{\bm{v}}_{O})$. Under this framework, the log-likelihood $p_{w}({\bm{v}}_{O})$ can be optimized using an EM algorithm: during the E-step, we fix $w$ and update the variational distribution $q$; during the M-step, we update $w$ to maximize the log-likelihood of all the entity pairs, i.e., $E_{q({\bm{v}}_{H})}[\log p_{w}({\bm{v}}_{O},{\bm{v}}_{H})]$, as illustrated in Figure 1.
@@ 3.2 E-step: inference
In this step, we fix $w$ and update $q_{\theta}$ to minimize the KL divergence $D_{KL}$. Directly minimizing $D_{KL}$ is intractable, as it involves computing the entropy of $q_{\theta}$. Therefore, we follow Qu & Tang (2019) and optimize the reverse KL divergence of $q_{\theta}$ and $p_{w}$, leading to the following objective:
@@ 4.1 Efficient optimization via logical deduction
Inference and learning with logic rules of length $L$ can be computationally intensive, as the search space for paths grows exponentially with increasing $L$. To enhance reasoning efficiency, we decompose a rule in equation 1 using logic deduction, inspired by Cheng et al. (2023) in KG completion:
@@
Here $\bigwedge_{k=1}^{L}r_{k}(e_{k-1},e_{k})$ represents the path formed by $r_{1},r_{2},...,r_{L}$ connecting $e_{i}$ to $e_{j}$ with $e_{0}=e_{i}$ and $e_{k}=e_{j}$. This can be reorganized as a series of single-step logic reasoning:
@@
In this way, each logic rule of length $L$ can be viewed as a deductive combination of $L$ short rules of length 1. At each step, following Suchanek et al. (2012), we perform one-step inference to update $p_{w}({\bm{v}}_{(e,e^{\prime})})$ for each $(e,e^{\prime})\in{\mathcal{H}}$ by aggregating the alignment probability from neighbors:
$1-\prod_{\begin{subarray}{c}(e,r,e_{t})\in\mathcal{T},\\ (e^{\prime},r^{\prime},e_{t}^{\prime})\in\mathcal{T}^{\prime}\end{subarray}}\left(1-\eta(r)p_{sub}(r\subseteq r^{\prime})p_{w}({\bm{v}}_{(e_{t},e_{t}^{\prime})})\right)\times\left(1-\eta(r^{\prime})p_{sub}(r^{\prime}\subseteq r)p_{w}({\bm{v}}_{(e_{t},e_{t}^{\prime})})\right).$ (9)
where $\eta(r)$ is a relation pattern of $r$ measuring the uniqueness of $e$ through relation $r$ given a specified tail entity $e_{t}$, quantified by $\eta(r)=\frac{|\{e_{t}|(e_{h},r,e_{t})\in{\mathcal{T}}\}}{|\{(e_{h},e_{t})|(e_{h},r,e_{t})\in{\mathcal{T}})\}|}$. $p_{sub}(r\subseteq r^{\prime})$ denotes the probability that relation $r$ is a subrelation of $r^{\prime}$. This formulation enables inference with confidence by explicitly quantifying confidence $w$ using $\eta$ and $p_{sub}(r\subseteq r^{\prime})$. In this way, the update of the weight $w$ simplifies to updating $p_{sub}(r\subseteq r^{\prime})$ during the M-step (equation 6), as $\eta(r)$ for each relation $r$ is constant. In practice, the update of $p_{sub}(r\subseteq r^{\prime})$ can be computed by:
$\frac{\sum\left(1-\prod_{(e_{h}^{\prime},r^{\prime},e_{t}^{\prime})\in{\mathcal{T}}^{\prime}}\left(1-{\bm{v}}_{(e_{h},e_{h}^{\prime})}{\bm{v}}_{(e_{t},e_{t}^{\prime})}\right)\right)}{\sum\left(1-\prod_{e_{h}^{\prime},e_{t}^{\prime}\in{\mathcal{E}}^{\prime}}\left(1-{\bm{v}}_{(e_{h},e_{h}^{\prime})}{\bm{v}}_{(e_{t},e_{t}^{\prime})}\right)\right)}.$ (10)

### v2

# NeuSymEA: Neuro-symbolic Entity Alignment via Variational Inference

@@ Abstract
Entity alignment (EA) aims to merge two knowledge graphs (KGs) by identifying equivalent entity pairs. Existing methods can be categorized into symbolic and neural models. Symbolic models, while precise, struggle with substructure heterogeneity and sparsity, whereas neural models, although effective, generally lack interpretability and cannot handle uncertainty. We propose NeuSymEA, a unified neuro-symbolic reasoning framework that combines the strengths of both methods to fully exploit the cross-KG structural pattern for robust entity alignment. NeuSymEA models the joint probability of all possible pairs’ truth scores in a Markov random field, regulated by a set of rules, and optimizes it with the variational EM algorithm. In the E-step, a neural model parameterizes the truth score distributions and infers missing alignments. In the M-step, the rule weights are updated based on the observed and inferred alignments, handling uncertainty. We introduce an efficient symbolic inference engine driven by logic deduction, enabling reasoning with extended rule lengths. NeuSymEA achieves a significant 7.6% hit@1 improvement on $\text{DBP15K}_{\text{ZH-EN}}$ compared with strong baselines and demonstrates robustness in low-resource settings, achieving 73.7% hit@1 accuracy on $\text{DBP15K}_{\text{FR-EN}}$ with only 1% pairs as seed alignments. Codes are released at https://github.com/chensyCN/NeuSymEA-NeurIPS25.
@@ 1 Introduction
Knowledge graphs (KGs) are crucial for organizing structured knowledge about entities and their relationships, enhancing search capabilities across various applications. They are widely used in question-answering systems [1, 2], recommendation systems [3, 4], SQL generation [5, 6], natural language processing [7, 8, 9], etc.. Despite their utility, real-world KGs often face issues like incompleteness, domain specificity, and language constraints, which hinder their effectiveness in cross-disciplinary or multilingual contexts [10, 11]. To address these issues, entity alignment (EA) aims to merge disparate KGs into a unified, comprehensive knowledge base by identifying and linking equivalent entities across different KGs. For example, aligning entities between a biomedical KG and a pharmaceutical KG allows for mining cross-discipline relationships through the aligned entities, such as identifying the same drugs and their effects on different diseases to enhance drug repurposing efforts. This alignment enables more nuanced exploration and interrogation of interconnected data, providing richer insights into how entities function across multiple domains.
@@
Entity alignment models aim to determine the equivalence of two entities by assessing their alignment probability. Existing methods can be broadly categorized into symbolic models and neural models. Symbolic models [12, 13, 14] provide interpretable and precise inference by mining ground rules, but they struggle with aligning low-degree entities, especially those without aligned neighbors. In such cases, the lack of supporting rules leads to low recall. Conversely, neural models, such as translation models [15, 16] and graph convolutional networks (GCNs) [17, 18, 19, 20, 21, 22, 23], excel in recalling similar entities by embedding them in a continuous space, yet they often fail to distinguish entities with similar representations, causing a drop in precision as the entity pool grows. Neuro-symbolic models aim to combine the strengths of both approaches, offering robust reasoning ability for entity alignment in challenging scenarios.
@@
However, neuro-symbolic reasoning in entity alignment (EA) faces several challenges. First, combining symbolic and neural models into a unified framework is suboptimal due to the difficulty in aligning their objectives. Current approaches either use neural models as auxiliary modules for symbolic models to measure entity similarity [14] or employ symbolic models to refine pseudo-labels [24, 25]. Second, in EA task, the search space for rules is large. Deriving ground rules from both intra-KG and inter-KG structural patterns results in an exponentially growing search space as rule length increases, complicating efficient rule weight estimation and inference. Finally, generating interpretations for EA remains underexplored. Effective interpretations should not only generate supporting rules but also quantify their confidence through rule weights.
@@
To overcome these challenges, we propose NeuSymEA, a neuro-symbolic framework that combines the strengths of both symbolic and neural models. NeuSymEA models the joint probability of truth score assignment for all possible entity pairs using a Markov random field, regulated by a set of weighted rules. This joint probability is optimized via a variational EM algorithm. During the E-step, a neural model parameterizes the truth scores and infers the missing alignments. In the M-step, the rule weights are updated based on both observed and inferred alignments. To leverage long rules without suffering from the exponential search space, we employ logic deduction to decompose rules of any length into a set of unit-length sub-rules. This allows for efficient inference and weight updates for long rules. After training, the missing alignments are jointly inferred by both components. Additionally, we introduce an explainer to enhance interpretability. By reversing the rule decomposition process, we extract long rules as explicit supporting evidence for alignments and recover rule weights as quantified confidence scores. Our contributions are summarized as below:
  • A principled neuro-symbolic reasoning framework via variational EM: While variational EM has been utilized in KG completion tasks [26, 27], adapting it directly to the EA task is nontrivial because they only consider single-KG structures. We bridge this gap by formulating truth scores and weighted cross-KG rules, and modeling the joint probability of the truth scores in a Markov random field regulated by the weighted cross-KG rules.
@@
  • Empirical validation and superior results: NeuSymEA demonstrates state-of-the-art performance on benchmark datasets, delivering robust alignment accuracy and rule-based interpretations.
@@ 2.1 Problem statement
A knowledge graph $\mathcal{G}$} comprises a set of entities $\mathcal{E}$}, a set of relations $\mathcal{R}$}, and a set of relation triples $\mathcal{T}$} where each triple $(e_{i},r_{k},e_{j})\in{\mathcal{T}}$ represents a directional relationship between its head entity and tail entity. Given two KGs ${\mathcal{G}}=\{{\mathcal{E}},{\mathcal{R}},{\mathcal{T}}\}$, ${\mathcal{G}}^{\prime}=\{{\mathcal{E}}^{\prime},{\mathcal{R}}^{\prime},{\mathcal{T}}^{\prime}\}$, and a set of observed aligned entity pairs ${\mathcal{O}}=\{(e_{i},e_{i}^{\prime})|e_{i}\in{\mathcal{E}},e^{\prime}_{i}\in{\mathcal{E}}^{\prime}\}_{i=1}^{n}$, the goal of entity alignment is to infer the missing alignments by reasoning with the existing alignments. This problem can be formulated in a probabilistic way: each pair $(e,e^{\prime}),e\in{\mathcal{E}},e^{\prime}\in{\mathcal{E}}^{\prime}$ is associated with a binary indicator variable ${\bm{v}}_{(e,e^{\prime})}$. ${\bm{v}}_{(e,e^{\prime})}=1$ means $(e,e^{\prime})$ is an aligned pair, and ${\bm{v}}_{(e,e^{\prime})}=0$ otherwise. Given some observed alignments ${\bm{v}}_{O}=\{{\bm{v}}_{(e,e^{\prime})}=1\}_{(e,e^{\prime})\in{\mathcal{O}}}$, we aim to predict the labels of the remaining hidden entity pairs ${\mathcal{H}}={\mathcal{E}}\times{\mathcal{E}}^{\prime}\backslash{\mathcal{O}}$, i.e., ${\bm{v}}_{H}=\{{\bm{v}}_{(e,e^{\prime})}\}_{(e,e^{\prime})\in{\mathcal{H}}}$.
@@ 2.2 Symbolic reasoning for entity alignment
$w_{p,p^{\prime}}:(e_{j}\equiv e_{j}^{\prime})\wedge p(e_{i},e_{j})\wedge p^{\prime}(e_{i}^{\prime},e_{j}^{\prime})\Longrightarrow(e_{i}\equiv e_{i}^{\prime}),$ (1)
@@ 3 Neuro-symbolic reasoning framework for entity alignment
Figure 1: Framework illustration of NeuSymEA. The yellow solid line represents a pre-aligned entity pair. The symbolic model computes the path-level matching probability of entity pairs by mining rules with learned weights. The neural model learns embeddings and calculates entity-level matching scores. NeuSymEA models their agreement using a joint probability distribution over observed pairs and parameterized truth scores for hidden pairs, optimizing through a variational EM algorithm.
@@ 3.1 Variational EM
here, $q({\bm{v}}_{H})$ is a variational distribution of the hidden variables ${\bm{v}}_{H}$. This inequality holds for all $q$ because $p_{w}({\bm{v}}_{O})=\text{ELBO}(q,{\bm{v}}_{O};w)+D_{KL}(q({\bm{v}}_{H})\|p_{w}({\bm{v}}_{H}\mid{\bm{v}}_{O}))$, where $D_{KL}(q({\bm{v}}_{H})\|p_{w}({\bm{v}}_{H}\mid{\bm{v}}_{O}))\geq 0$ is the KL-divergence between $q({\bm{v}}_{H})$ and $p_{w}({\bm{v}}_{H}\mid{\bm{v}}_{O})$. Under this framework, the log-likelihood $p_{w}({\bm{v}}_{O})$ can be optimized using an EM algorithm, an efficient method to find the maximum likelihood where the model depends on unobserved hidden variables: during the E-step, we fix $w$ and update the variational distribution $q$; during the M-step, we update $w$ to maximize the log-likelihood of all the entity pairs, i.e., $E_{q({\bm{v}}_{H})}[\log p_{w}({\bm{v}}_{O},{\bm{v}}_{H})]$, as illustrated in Figure˜1.
@@ 3.2 E-step: inference
In this step, we fix $w$ and update $q_{\theta}$ to minimize the KL divergence $D_{KL}$. Directly minimizing the KL divergence is intractable, as it involves computing the entropy of $q_{\theta}$. Therefore, we follow [26] and optimize the reverse KL divergence of $q_{\theta}$ and $p_{w}$, leading to the following objective:
@@ 4.1 Efficient optimization via logical deduction
Inference and learning with logic rules of length $L$ can be intractable, as the search space for paths grows exponentially with increasing $L$. To enhance reasoning efficiency, we decompose a rule in Equation 1 using logic deduction, inspired by [28] in KG completion:
@@
Here $\bigwedge_{k=1}^{L}r_{k}(e_{k-1},e_{k})$ represents the path formed by $r_{1},r_{2},...,r_{L}$ connecting $e_{i}$ to $e_{j}$ with $e_{0}=e_{i}$ and $e_{k}=e_{j}$. This long rule can be reorganized as the combination of a series of unit-length logic reasoning:
@@
In this way, each logic rule of length $L$ can be viewed as a deductive combination of $L$ short rules of length 1. At each step, following [12], we perform one-step inference to update $p_{w}({\bm{v}}_{(e,e^{\prime})})$ for each $(e,e^{\prime})\in{\mathcal{H}}$ by aggregating the alignment probability from neighbors:
$1-\prod_{\begin{subarray}{c}(e,r,e_{t})\in\mathcal{T},\\ (e^{\prime},r^{\prime},e_{t}^{\prime})\in\mathcal{T}^{\prime}\end{subarray}}\left(1-\eta(r)p_{sub}(r\subseteq r^{\prime})p_{w}({\bm{v}}_{(e_{t},e_{t}^{\prime})})\right)\times\left(1-\eta(r^{\prime})p_{sub}(r^{\prime}\subseteq r)p_{w}({\bm{v}}_{(e_{t},e_{t}^{\prime})})\right),$ (9)
where $\eta(r)$ is a relation pattern of $r$ measuring the uniqueness of $e$ through relation $r$ given a specified tail entity $e_{t}$, quantified by $\eta(r)=\frac{|\{e_{t}|(e_{h},r,e_{t})\in{\mathcal{T}}\}}{|\{(e_{h},e_{t})|(e_{h},r,e_{t})\in{\mathcal{T}})\}|}$. $p_{sub}(r\subseteq r^{\prime})$ denotes the probability that relation $r$ is a subrelation of $r^{\prime}$. This technique enables inference with confidence by explicitly quantifying confidence $w$ during each inference step by introducing $\eta$ and $p_{sub}(r\subseteq r^{\prime})$. Moreover, in this way, the update of the weight $w$ simplifies to updating $p_{sub}(r\subseteq r^{\prime})$ during the M-step (Equation 6), as $\eta(r)$ for each relation $r$ is constant. In practice, the update of $p_{sub}(r\subseteq r^{\prime})$ can be computed by:
$\frac{\sum\left(1-\prod_{(e_{h}^{\prime},r^{\prime},e_{t}^{\prime})\in{\mathcal{T}}^{\prime}}\left(1-{\bm{v}}_{(e_{h},e_{h}^{\prime})}{\bm{v}}_{(e_{t},e_{t}^{\prime})}\right)\right)}{\sum\left(1-\prod_{e_{h}^{\prime},e_{t}^{\prime}\in{\mathcal{E}}^{\prime}}\left(1-{\bm{v}}_{(e_{h},e_{h}^{\prime})}{\bm{v}}_{(e_{t},e_{t}^{\prime})}\right)\right)},$ (10)
@@
In Section˜A.2, we provided a detailed complexity analysis, demonstrating that parameter complexity scales linearly with dataset size, while computational complexity is quadratic. Our implementation enables efficient execution through parallel computing and batch processing.